Circular No. 8578

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2002 EX_{12} (NEAT) An apparently asteroidal object reported by NEAT, and designated

2002 EX_{12} (discovery observation given below; cf. MPS 52825), has been reported as showing a cometary appearance by observers at two different sites. B. D. Warner, Colorado Springs, CO, writes that his CCD frames taken with a 0.35-m Schmidt-Cassegrain reflector on 2005 July 28.26 UT show a 77′′ tail in p.a. 144^{o}, while frames taken on July 29.26 show a 90′′ tail in p.a. 145^{o}. A. Fitzsimmons, Queen’s University, Belfast, reports that the object showed no coma but a faint, straight tail extending 30′′ in p.a. 145^{o} on R-band images taken on July 29.5 with the 2.0-m ‘Faulkes Telescope North’ at Haleakala; he adds that observations with that telescope on 2005 May 10.5 and with the 3.5-m New Technology Telescope on May 14.3 did not obviously show a tail.

2002 UT  α_{2000}  δ_{2000}  Mag.
Mar. 15.26990  12^{h}19^{m}20.^{s}62  −0^{o}28′06′′.9  19.8

The following orbital elements are taken from MPEC 2005-P01:

Epoch = 2005 Aug. 18.0 TT

T = 2005 Sept.17.8587 TT  ω = 217.^{o}9272
e = 0.767520  Ω = 176.2462 }2000.0
q = 0.605325 AU  i = 11.3189

a = 2.603775 AU n^{o} = 0.2345843 P = 4.202 years

COMET C/2005 O1 (NEAT) Another apparently asteroidal object found by NEAT (discovery ob-

servation below) has also been found to show cometary activity by Fitzsimmons, his R-band CCD frames taken with the 2.0-m reflector on July 31.5 UT showing a roughly symmetrical coma extending to 3′′.5 from the object’s condensation.

2005 UT  α_{2000}  δ_{2000}  Mag.
July 27.39588  22^{h}31^{m}49.^{s}38  −21^{o}25′10′′.7  19.1

The available astrometry, preliminary parabolic orbital elements (T = 2005 May 23.987 TT, q = 3.61757 AU, ω = 325^{o}.726, Ω = 304^{o}.594, i = 156^{o}.196, equinox 2000.0), and an ephemeris appear on MPEC 2005-P09.

2005 August 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green